# Eagle: Efficient Training-Free Router for Multi-LLM Inference

### v1

@@ 1 Introduction
Existing approaches to LLM routing have limitations. Routerbench [10] proposes using traditional machine-learning methods like MLP and KNN to predict quality rankings, but these require heavy retraining and cannot utilize real-time user feedback. RouteLLM [17] suggests a similarity-weighted ranking method for training-free prediction, but our experiments indicate that relying solely on partial information can lead to biased results and suboptimal overall performance.
@@ 2.1 Eagle Design
2. Efficient Feedback Integration Since user feedback typically only provides a preference between two models, obtaining a complete ranking of all models from users is challenging. To address this, we employ the ELO ranking method, which transforms sparse pairwise comparisons into a comprehensive ranking system, maximizing the value of each user interaction.

### v2

@@ 1 Introduction
Existing approaches to LLM routing have limitations. Routerbench [11] proposes using traditional machine-learning methods like MLP and KNN to predict quality rankings, but these require heavy retraining and cannot utilize real-time user feedback. RouteLLM [18] proposes a similarity-weighted ranking approach for training-free prediction. While effective for binary routing decisions, their method is limited to scenarios involving only two models and does not address cases with $n\geq 3$ models (where $n$ denotes the number of models).
@@ 2.1 Eagle Design
2. Efficient Feedback Integration Since user feedback typically only provides a preference between two models, obtaining a complete ranking of all models from users is challenging. To address this, we employ the ELO[9] ranking method, which transforms sparse pairwise comparisons into a comprehensive ranking system, maximizing the value of each user interaction.